OpenStudy (anonymous):

if f(x)=7x and g(x)=3x+1, find (fog)(x)

OpenStudy (i_need_help_please):

g (x) = 7x +1 fog (x) = 3(7x+1) fog(x) = 21x +3 fog(4) = 84+3 fog(x) = 87

OpenStudy (anonymous):

Thanks